Grade 126 grade maths

The distance between the school and a student’s house is 1 km 875 m. Everyday she walks both ways. Find the total distance covered by her in six days.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Distance covered between school and house = 1 km 875 m = 1000 + 875 = 1875 m
Since, the student walk both ways.
Hence, distance travelled by the student in one day = 2×1875 = 3750 m
Distance travelled by the student in 6 days = 3750 m×6 = 22500 m = 22 km 500 m∴Total distance covered by the student in six days is 22 km and 500 m

Thanks